Assume Party Store has the following account balances for the month of March 2021, and that the company uses a perpetual inventory system.

  

Sales revenue                           $75,800         Cost of goods sold                  $38,500

                                                                      Inventory (Mar. 31, 2021)           1,800

Advertising expense                    5,200         Insurance expense                       1,700

Rent expense                               3,300         Sales discounts                           2,900

Gain on sale of building              6,900         Salaries expense                          8,200

Inventory (Mar. 1, 2021)              2,200         Income tax expense                    6,100

Required:

1.   Prepare a multiple-step income statement for the month ended March 31, 2021.

2.   Calculate the inventory turnover ratio for the month of March.

      Would you expect this ratio to be higher or lower in December 2021? Explain.

3.   Calculate the gross profit ratio for the month of March.

Answer:  

Requirement 1

                                 Party Store

                 Multiple-Step Income Statement

            For the month ended March 31, 2021

Net sales:                                                          

Total sales revenue                   $75,800           

Net sales                                                           

                                                                        

Gross Profit                                                       

                                                                        

Total                                                                18,400

Operating income                                             16,000

Nonoperating items:                                         

                                                                        

Income before income taxes                             

                                                                        

Net income                                                      $16,800
Answer #1
Multiple Income Statement
Particulars Amount ($) Amount ($)
Sales Revnue 75800
Less: Sales Discount 2900
Net Sales 72900
Less: Cost of goods sold $38,500.00
Gross Profit $34,400.00
Less: Operating Expenses
Advertising Expense $5,200.00
Rent Expense $3,300.00
Salaries Expense $8,200.00
Insurance expense $1,700.00
Total Operating expenses $18,400.00
Operating Income (a) $16,000.00
Non Operating Income/ (Loss):
Gain on Sale of Building $6,900.00
Total non operating income (b) $6,900.00
Income before tax (c=a+b) $6,100.00
Less: Tax expense $1,830.00
Net Income $4,270.00
Inventory Turnover Ratio= Cost of Goods Sold/ Average Inventory
38500/((2200+1800)/2)= 9.625
Inventory Turnover ratio is higher the better and expecting to more higher turnover ratio in dec 2021 because higher turnover ratio is better
Gross Profit Ratio= Gross profit/ Net Sales
=34400/72900= 47.19%